OpenStudy (anonymous):

write the solution in lowest terms -1/3 + 3/4

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\] \[-\frac{1}{3}+\frac{3}{4}=\frac{-1\times 4+3\times 3}{3\times 4}=\frac{-4+9}{12}=\frac{5}{12}\]

OpenStudy (anonymous):

thx